# BiM-VFI: Bidirectional Motion Field-Guided Frame Interpolation for Video with Non-uniform Motions

### v2

@@ 1 Introduction
The contributions of our work are summarized as:
  • A novel motion description map, Bidirectional Motion fields (BiM) is proposed to describe non-uniform motions effectively;
  • A Knowledge Distillation for VFI-Centric Flow supervision (KDVCF) is presented to coincide the two purposes of direct supervision on optical flow estimation and photometric reconstruction during training;
  • A BiM-guided FlowNet (BiMFN), with Content-Aware Upsampling Network (CAUN) for precise optical flow estimation.
  • Our BiM-VFI model significantly surpasses the very recent state-of-the-art VFI methods by 26% and 45% improvements in LPIPS [43] and STLPIPS [5] respectively, yielding interpolated frames with much less blurs at arbitrary target time instances.
@@ 3.1 Bidirectional Motion Field (BiM)
where $H$ and $W$ are the height and width of desired bidirectional optical flows, $\textbf{V}_{t\rightarrow 0}$ and $\textbf{V}_{t\rightarrow 0}$. We demonstrate that our BiM, under the uniform motion assumption, can interpolate frames with a similar sense of time indexing as described in Sec. 4.5.
@@ 4.1 Experiments details
Implementation details. We trained our BiM-VFI with a training split of Vimeo90k septuplet datasets [41]. We randomly crop the images to a resolution of 256 $\times$ 256, flip horizontally and vertically, rotate, reverse temporally, and permute the color channels to augment the training data. We set the batch size to 32, and train the model for 400 epochs with an initial learning rate of 1 ×$\times 10^{-4}$. We gradually decay the initial learning rate using a Cosine annealing scheduler [21] and optimize our model using the AdamW optimizer [20]. Also, because the architecture of our BiM-VFI is based on a recurrent pyramid architecture, we employed resolution-aware adaptation for the pyramid hierarchy depth proposed by Jin et al. [13]. For training on Vimeo90K, we used 3 pyramid levels, while 5 pyramid levels are used for SNU-FILM [3] and SNU-FILM-arb [6], and 7 pyramid levels for Xtest [35]. As mentioned, our proposed KDVCF computes the BiM during training, and for inference time, the BiM is represented according to Eq. 2 corresponding to a uniform motion scenario.
@@ 4.4 Ablation studies
4.5 Uniform Motion Assumption
Our BiM-VFI is limited in interpolating frames with uniform motion at inference time, due to the unavailability of the target BiM. However, Fig. 6 demonstrates that other VFI methods, such as those employing time indexing (EMA-VFI [45]) and distance indexing ([D,R]-EMA-VFI [44]), also fail to adequately interpolate the target frame with non-uniform motions, where the boundary of the tree at the target frame (indicated by the blue line) does not align with the interpolation results from EMA-VFI, [D,R]-EMA-VFI, or our BiM-VFI (indicated by the green line). Moreover, the boundary of the tree from the interpolated frame using EMA-VFI, which employs time indexing, is well aligned with those of models using distance indexing or BiM under the uniform motion assumption. The results above suggest that time indexing implicitly applies the uniform motion assumption in a data-driven manner, thereby supporting that the use of uniform motion assumption at inference is valid.
@@ 5 Conclusion
[…]
@@ B.1 Quantitative Results
In Tab. 5, we also provided pixel-centric metrics measured on fixed-time datasets (Vimeo 90K-triplet [41], SNU-FILM [3], and XTest [35] single) and complexity comparisons between other SOTA methods. We measured FLOPs for interpolating 1280$\times$720-sized source images and the total parameters used in the methods. As shown in Tab. 5, our BiM-VFI effectively reduced the number of parameters by employing a recurrent pyramid architecture, while having moderate computational complexity among the other SOTA methods in terms of FLOPs.

### v4

@@ 3.1 Bidirectional Motion Field (BiM)
where $H$ and $W$ are the height and width of desired bidirectional optical flows, $\textbf{V}_{t\rightarrow 0}$ and $\textbf{V}_{t\rightarrow 1}$. We demonstrate that our BiM, under the uniform motion assumption, can interpolate frames with a similar sense of time indexing as described in Sec. 4.5.
@@ 4.4 Ablation studies
4.5 Limitation on Uniform Motion Assumption
Our BiM-VFI is limited for frame interpolation under a uniform motion assumption at inference time, due to the unavailability of the target BiM, which is an inherit limitation as for all other VFI methods. Fig. 6 demonstrates that other VFI methods, such as those employing time indexing (EMA-VFI [45]) and distance indexing ([D,R]-EMA-VFI [44]), also fail to adequately interpolate the target frame with non-uniform motions, where the boundary of the tree at the target frame (indicated by the blue line) does not align with the interpolation results from EMA-VFI, [D,R]-EMA-VFI, or our BiM-VFI (indicated by the green line). Moreover, the boundary of the tree from the interpolated frame using EMA-VFI that employs time indexing is well aligned with those of models using distance indexing or BiM under the uniform motion assumption. This suggests that the time-indexing-based method implicitly tends to comply with the uniform motion assumption from training where uniform motion is dominant, thereby supporting that the uniform motion assumption at inference is more likely enforced for all the methods in VFI.
@@ 5 Conclusion
6 Acknowledgement
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government [Ministry of Science and ICT (Information and Communications Technology)] (Project Number: RS-2022-00144444, Project Title: Deep Learning Based Visual Representational Learning and Rendering of Static and Dynamic Scenes, 100%).
@@ 6 Acknowledgement
In this supplementary material, we first provide additional details of our proposed BiM-VFI. Especially, the detailed network structure of CAUN and SN, loss functions, implementation details, and the proof of how BiM can describe bidirectional motion distinctly are explained in Appendix A. Subsequently, in Appendix B, we provided additional experimental results that could not be included in the main paper due to the page limitation. In Sec. B.1, pixel-centric metrics (PSNR and SSIM) in SNU-FILM-arb [6], Vimeo90K-triplet [41], SNU-FILM [3], and XTest-single [35], and ×$\times 8$ interpolation on XTest dataset are provided. Also, in Sec. B.4, we provided the results of the user study we conducted for interpolated videos from various VFI methods. Lastly, in Sec. B.5, we provided additional qualitative comparisons on SNU-FILM-arb [6] datasets.
@@ A.3 Loss Functions
A.4 Implementation details
We trained our BiM-VFI with a training split of Vimeo90k septuplet datasets [41]. We randomly crop the images to a resolution of 256 $\times$ 256, flip horizontally and vertically, rotate, reverse temporally, and permute the color channels to augment the training data. We set the batch size to 32, and train the model for 400 epochs with an initial learning rate of 1 ×$\times 10^{-4}$. We gradually decay the initial learning rate using a Cosine annealing scheduler [21] and optimize our model using the AdamW optimizer [20]. Also, because the architecture of our BiM-VFI is based on a recurrent pyramid architecture, we employed resolution-aware adaptation for the pyramid hierarchy depth proposed by Jin et al. [13]. For training on Vimeo90K, we used 3 pyramid levels, while 5 pyramid levels are used for SNU-FILM [3] and SNU-FILM-arb [6], and 7 pyramid levels for Xtest [35]. As mentioned, our proposed KDVCF computes the BiM during training, and for inference time, the BiM is represented according to Eq. 2 corresponding to a uniform motion scenario.
@@ B.1 Quantitative Results
In Tab. 5, we also provided pixel-centric metrics measured on fixed-time datasets (Vimeo 90K-triplet [41], SNU-FILM [3], and XTest [35] single) and complexity comparisons between other SOTA methods.
B.2 Computational complexity
We additionally provide below #’s of parameters and FLOPs on each component for 256$\times$256-sized images.
MFE CFE BiMFN CAUN SN Total
#Params(M) 0.58 0.58 3.41 0.61 1.7 6.88
FLOPs(G) 12.02 12.02 18.81 11 24.64 78.49
We measured FLOPs for interpolating 1280$\times$720-sized source images and the total parameters used in the methods. As shown in Tab. 5, our BiM-VFI effectively reduced the number of parameters by employing a recurrent pyramid architecture, while having moderate computational complexity among the other SOTA methods in terms of FLOPs.
GIMM-VFI-R EMA-VFI UPR AMT Ours
#Params(M) 19.73 65.66 6.56 30.64 6.88
FLOPs(G) 9187 1714 1228 2395 1177
Runtime(ms) 494 104 53 183 151
B.3 Additional ablation study
We provide below the detection performance of small objects and object boundaries with and without CAUN module. As shown, the CAUN can help capture well small toes (top) and detect tight boundaries of the windmill blade (bottom), while failing without it.
Also, as mentioned in Suppl., while our KDVCF increases the training time from 2.5 to 4 days using 4 A100 GPUs due to the $\mathcal{P}_{\mathcal{T}}$ process, it does not increase the inference time.
We also compared the #’s of parameters, FLOPs, and runtime (measured on an A100 GPU with 1280$\times$768-sized images) in the below table. It can be noted that our BiM-VFI has a low number of parameters, showing moderate runtime.